A bus went 318 km farther than a car.The car went one third as far as the bus.How far did each vehicle travel.
Choices 147,151,159

Answer by edjones(7311)

let c=car's distance b=bus distance
b=c+318
3c=b
3c=c+318
subtract c from each side: 2c=318
divide 2 into each side: c=159 km
b=3*159=477 km
Check:
477-159=318
Ed